OpenStudy (anonymous):

Identify the 31st term of an arithmetic sequence where a1 = 26 and a22 = –226.

OpenStudy (anonymous):

Well first find d. a1 = 26 a22 = –226 a22=a1+21d Substitute! -226=26+21d 21d=-252 d= -12 Now lets find the 31st term. a31=a1+30d a31=26+30 * (-12) a31=26 + -360 a31= -334

OpenStudy (anonymous):

how did you get the d part

OpenStudy (anonymous):

\[t_{n}=a+(n-1)d\]

OpenStudy (anonymous):

what?.. i don't get it

OpenStudy (anonymous):

\[t_{n}\] is the nth term.In this case I put it as 22nd term.d is the common diff. which was found.And a is a1

OpenStudy (anonymous):

so its always lower like if it was 17 it would be a1+16?

OpenStudy (anonymous):

a1+16d

OpenStudy (anonymous):

okay let me see if i get this one by myself

OpenStudy (anonymous):

a1 = –12 and a27 = 66. a1=-12 a27=66 a27=a1+26d?